# JFR: A Jump Frontier Relaxation Strategy for Fast Bellman–Ford

### v2

@@ 4.4 C++ Experiments: Large-Scale Randomized Benchmarking
To validate JFR in large-scale scenarios, we conducted extensive randomized benchmarking across sparse and dense graphs. Each graph was repeatedly generated and tested to obtain stable averages:
@@
Table 2: C++ Experiments: JFR vs SPFA-SLF (Runtime and Relaxation Ops, Averaged over Large-Scale Tests)
@@
Table 3: Nonlinear Acceleration Validation (Original Graphs, No Edge Increment)
Instance —V— JFR Time [ms] JFR Ops NegDense_XL_1 4951 606.43 19,900,385 NegDense_XL_2 4538 1175.97 27,425,775
a baseline for comparison
Table 4: Nonlinear Acceleration Validation (After Edge Increment)
Instance —V— Edges (B) JFR Time B [ms] JFR Ops B NegDense_XL_1 4951 4,922,884 449.48 9,936,365 NegDense_XL_2 4538 5,682,231 710.59 13,797,216
Nonlinear Acceleration Phenomenon. Across both NegDense_XL instances, a small edge increment (approximately +8.8% and +9.6%) unexpectedly causes both JFR runtime and relaxation operations to decrease—sometimes by 25–50%. This counterintuitive behavior reveals a nonlinear acceleration effect intrinsic to JFR: when the vertex set is fixed, additional edges can shift the graph into a more connectivity-rich regime where Frontier Filtering becomes more aggressive and jump propagation stabilizes earlier. As a result, multiple Bellman–Ford iterations collapse into fewer Bounded Local Propagation Steps, sharply reducing redundant relaxations and total work.
@@ 4.6 Scalability and Extensibility Analysis
Table 5: Scalability Benchmark: JFR vs SPFA-SLF
@@
NegDense_Ultra-1 SPFA-SLF 14,893.03 566,687,820 PASS
JFR 12,636.77 136,744,212 PASS
NegDense_Ultra-2 SPFA-SLF 25,199.07 1,188,649,749 PASS
JFR 23,663.12 547,254,897 PASS
Table 6: Performance Comparison on Large-Scale Adversarial Graph ($N=500,000$)

### v3

@@ 4.4 C++ Experiments: Large-Scale Randomized Benchmarking
To validate JFR in large-scale scenarios, we conducted extensive randomized benchmarking across sparse and dense graphs. We performed comparative evaluations using both a standard default environment and an aggressive -O3 optimized environment, distinguishing between algorithmic logic overhead and implementation-level efficiency. Each graph was repeatedly generated and tested to obtain stable averages.
@@
Table 2: C++ Experiments (default): JFR vs SPFA-SLF (Runtime and Relaxation Ops, Averaged over Large-Scale Tests)
@@
Table 3: C++ Experiments (-O3 Optimized): JFR vs SPFA-SLF (Runtime and Relaxation Ops, Averaged over Large-Scale Tests)
Graph Algorithm Time (ms) Ops Check
Sparse_XL JFR 16.91 114,054 PASS
SPFA-SLF 11.18 1,490,188 PASS
NegDense_XL JFR 52.80 10,241,868 PASS
SPFA-SLF 56.63 20,721,610 PASS
Windmill_XL JFR 0.45 145,395 PASS
SPFA-SLF 0.28 195,611 PASS
SLF_Killer_XL JFR 12.51 1,014,790 PASS
SPFA-SLF 1,030.46 160,003,344 PASS
Impact of Compilation Environment
Comparing Table 2 (Default) and Table 3 (-O3 Optimized) reveals that JFR benefits disproportionately from compiler optimizations. In the default environment, the overhead of the std::priority_queue masks JFR’s algorithmic efficiency. However, under -O3 optimization, JFR not only drastically widens the performance gap on adversarial graphs (achieving an $\approx$82x speedup on SLF_Killer_XL) but also surpasses SPFA-SLF on dense negative graphs (NegDense_XL). This confirms that in realistic high-performance scenarios, the reduction in relaxation operations effectively amortizes the logarithmic queue overhead.
Table 4: Nonlinear Acceleration Validation (comparsion)
Instance —V— Edges (B) JFR Time B [ms] JFR Ops B NegDense_XL 4538 5,182,231 270.26 27,425,775 NegDense_XL 4538 5,682,231 190.55 13,797,216
Nonlinear Acceleration Phenomenon. Across the NegDense_XL instance, a small edge increment (approximately +9.6%) unexpectedly causes both JFR runtime and relaxation operations to decrease—sometimes by 25–50%. This counterintuitive behavior reveals a nonlinear acceleration effect intrinsic to JFR: when the vertex set is fixed, additional edges can shift the graph into a more connectivity-rich regime where Frontier Filtering becomes more aggressive and jump propagation stabilizes earlier. As a result, multiple Bellman–Ford iterations collapse into fewer Bounded Local Propagation Steps, sharply reducing redundant relaxations and total work.
@@ 4.6 Scalability and Extensibility Analysis
Table 5: Scalability Benchmark: JFR vs SPFA-SLF(-O3)
@@
NegDense_Ultra-1 SPFA-SLF 1947.27 327,064,005 PASS
JFR 383.51 68,869,589 PASS
NegDense_Ultra-2 SPFA-SLF 7771.65 1,188,649,749 PASS
JFR 7265.14 547,254,897 PASS
Table 6: Performance Comparison on Large-Scale Adversarial Graph ($N=500,000$,Default environment)